OpenStudy (anonymous):

HELP! I WILL GIVE MEDALS AND FAN! What is the slope of the equation 2y = 6x - 8?

OpenStudy (johnweldon1993):

Well you need this in the form y = mx + b *notice 'y' is by itself on that left side here, 'm' is the slope....and 'b' is the y-intercept... so with your equation 2y = 6x - 8 You need to get 'y' by itself...so you must divide the whole equation by 2 2y = 6x - 8 ----------- 2 2 2 y = 3x - 4 so compare that with the equation y = mx + b where 'm' is the slope y = mx + b y = 3x - 4 what is your slope?

OpenStudy (foreverandalways13):

3 because m=slope so whatever number is in "m" place thats youre slope